OpenStudy (anonymous):

can someone please explain to me how to solve this log problem :( 16^w+4^(w+1/2) = 3

OpenStudy (anonymous):

so far I have

OpenStudy (anonymous):

i'm stuck here

OpenStudy (raden):

i just use the exponent concept : 16^w + 4^(w+1/2) = 3 (4^w)^2 + 4^1/2 * 4^w - 3 = 0 (4^w)^2 + 2 * 4^w - 3 = 0 (4^w + 3)(4^w - 1) = 0 take each factor be equal zero, then solve for w

OpenStudy (anonymous):

ah ok i see... i have to get used to seeing the quadratic equation in the problems.

OpenStudy (raden):

yep, that i mean. so, did you get it ?

OpenStudy (anonymous):

yes w= 0 thanks =)

OpenStudy (anonymous):

from quadratic equation x= 1 so 4^w= 1 so w = 0

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

that's right, only w = 0 as the solution welcome :)